OpenStudy (anonymous):

What is the first step to solving the equation ? Choose the best answer. 1/3 x+18=0 A. Multiply both sides by 3. B. Add 18 to both sides. C. Subtract 18 from both sides. D. Divide both sides by 3

jhonyy9 (jhonyy9):

what is your opinion about C. ?

OpenStudy (mathmale):

This question is really for you, Mell. If left to your own resources, what's the first thing you'd do towards solving the equation for x?

OpenStudy (anonymous):

how would i add to both sides? 3*1/3 and 3*0? ?

OpenStudy (mathmale):

You want (1/3)x on the left side, all by itself. Subtract 18 from both sides and write your result here, please.

OpenStudy (anonymous):

idont know how to set it up

OpenStudy (mathmale):

Given: 1/3 x+18=0 Better written as (1/3) x+18=0 Subtract 18 from both sides: (1/3) x+18=0 -18=-18 --------------------- Finish this, please.

OpenStudy (anonymous):

i got 17.6 its higher than five so i can round up right?

OpenStudy (anonymous):

Move 18 to the RHS and then change it's sign. Divide each side by 1/3.

OpenStudy (danjs):

looks like the first step is to post on open study

OpenStudy (anonymous):

whats that supposed to mean?

OpenStudy (mathmale):

Please demonstrate how you obtained "i got 17.6." its higher than five so i can round up right? I asked you to subtract 18 from both sides of your equation. What is 18-18?

OpenStudy (danjs):

why cant you clear the fraction first..

OpenStudy (anonymous):

Dived each side by 1/3. Then move 54 to the RHS as -54.

OpenStudy (mathmale):

Mell: If you genuinely want to finish this discussion, please become involved in it.

OpenStudy (anonymous):

thanks for your help.